Determine the number of valence electrons in each of the followingelements.

Express your answer as an integer.

Ta

Hf

Fe

P
Answer #1
Ta 4f145d36s2 valenceelectron sum = 14 + 3 + 2 = 19

Hf 4f145d26s2 sum = 14 + 2+2 =18

Fe 3d64s2 sum = 6+2 =8

P 3s23p3  sum = 2 +3 =5